# Evaluating Grounded Reasoning by Code-Assisted LLMs for Mathematics

### v1

@@ Abstract
Assisting LLMs with code generation improved their performance on mathematical reasoning tasks. However, the evaluation of code-assisted LLMs is generally restricted to execution correctness, lacking a rigorous evaluation of their generated programs. In this work, we bridge this gap by conducting an in-depth analysis of code-assisted LLMs generated programs in response to math reasoning tasks. Our evaluation focuses on the extent to which LLMs ground their programs to math rules, and how that affects their end performance. For this purpose, we assess the generations of five different LLMs, on two different math datasets, both manually and automatically. Our results reveal that the distribution of grounding depends on LLMs’ capabilities and the difficulty of math problems. Furthermore, mathematical grounding is more effective for close-source models, while open-source models fail to employ math rules in their solutions correctly. On MATH500, the percentage of grounded programs decreased to half, while the ungrounded generations doubled in comparison to ASDiv grade-school problems. Our work highlights the need for in-depth evaluation beyond execution accuracy metrics, toward a better understanding of code-assisted LLMs capabilities and limits in the math domain.
Evaluating Grounded Reasoning by Code-Assisted LLMs for Mathematics
@@ 1 Introduction
Large Language Models (LLMs) have recently achieved outstanding performance on complex reasoning tasks such as mathematical reasoning, powered by scale and multi-step reasoning approaches, particularly Chain-of-Though (CoT)Wei et al. (2022), that require an LLM to generate the explicit reasoning steps, before generating the final answer.
Despite its success, investigating CoT reasoning steps revealed critical flows of LLMs, such as committing calculation errors Gao et al. (2023) and generating false positive chains Lyu et al. (2023), i.e: containing reasoning errors yet generating correct final answers.
@@
Code-assisted reasoning approaches Gao et al. (2023); Chen et al. (2022); Lyu et al. (2023); Gou et al. (2023); Mickus et al. (2024); Das et al. (2024) proposed to solve these problems by instructing LLMs to generate programmatic reasoning steps, e.g, Python programs, and delegate the execution of these programs to an interpreter, which avoids calculation errors and ensures faithfulness. Such approaches have been found to further improve LLMs performance on math reasoning tasks.
@@
However, performance improvement is predominantly evaluated by the correctness of the execution outcome Gao et al. (2023); Chen et al. (2022); Gou et al. (2023), rather than the quality of the generated programs and how it affects the reasoning process.
This is problematic, as the generated programs can be relaying on exhaustive searches or memorized information to generate correct answers instead of grounding the reasoning process to domain concepts, i.e: mathematical rules, producing less generalizable programs that are harder to verify and trust. Figure 1 show such programs generated by GPT-4 and StarCoder2 when solving math problems from the ASDiv data.
@@
Figure 1: Programs generated by GPT-4 (left) that uses a brute force search to find the answer to the input question, and StarCoder2 (right) that depends on memorized information rather than generating the steps to solve the input question.
@@
The goal addressed in this work is to evaluate the actual reasoning capabilities of code-assisted LLMs when solving mathematical tasks by analyzing their generated intermediate programs.
In this evaluation, we focus mainly on assessing the extent to which code-assisted LLMs ground their reasoning process to domain concepts and math rules when generating their solutions, and how grounding, or its absence, affects end performance on mathematical tasks.
Code enables access to domain concepts such as math rules, via libraries (e.g: Sympy), which can aid LLMs generate programs that are grounded to world knowledge (i.e: mathematical rules), aligning more to human-like reasoning. Given this we expect grounded programs to be trustworthy, easier to verify, and probably with a higher solve rate, while ungrounded programs to be more error-prone, harder to trust, and degrade end-performance.
We investigate the extent to which this is true by seeking answers to the following research questions.
  • RQ1: To what extent do code-assisted LLMs ground their reasoning to relevant mathematical concepts?
  • RQ2: How do code-assisted LLMs implement these concepts?
  • RQ3: What is the relationship between grounding to math concepts and end performance on mathematical tasks?
  • RQ4: How much do the answers to these questions vary with more challenging math problems?
@@
To answer these questions, we start by manually analyzing a subset of intermediate programs, generated by GPT4o-mini, GPT4, Qwen2.5, Llama3, and StarCoder2, for grounding and its implementation style, when solving math problems from the ASDiv and MATH500 datasets (3.3). Given the observations from the manual analysis, we design a taxonomy reflecting the different grounding levels used by evaluated models (3.4). To extend the analysis to the full set of generated programs, we employ two automated evaluation methods: an LLM-Judge using the latest o3-mini model, and our proposed method, Code-Structure Judge, that trains a Decision Tree classifier with features extracted from programs’ Abstract Syntax Tree (3.5). We find Code-Structure judge to achieve higher agreement with human judgment than LLM-judge, with $81\%$ and $73\%$, respectively (5.1); therefore, we employ it for the large-scale evaluation of all generated programs.
@@
To the best of our knowledge this is the first work to analyze generated programs of code-assisted LLMs on math reasoning tasks. We summarize our takeaways from Section (5.2) as following:
@@
    Code-assisted LLMs ground their programs in math knowledge either by invoking concepts from math library or implementing the concepts from scratch.
@@
    Code-assisted LLMs also relay on brute-force searches or memorized information to approach some math problems.
@@
    Grounded generations to math knowledge are more effective for GPT models and Qwen, while open-source models fail to correctly ground their reasoning to useful concepts.
@@
    The grounding distribution completely changed on the challenging MATH500 data, pushing GPT models and Qwen to double their reliance on brute-force search and memorized information, at the expense of more mathematically-grounded solutions.
Our in-depth evaluation highlights the need for more holistic assessment of LLMs generations, beyond execution accuracy metrics, which don’t provide insightful conclusions about models’ actual capabilities and limits in the math domain.
2 Related Work
Programs as Intermediate Steps for Math Reasoning Code-assisted reasoning approaches such as Chen et al. (2022); Gao et al. (2023); Imani et al. (2023); Lyu et al. (2023); Das et al. (2024) prompt LLMs to generate programs instead of intermediate steps in natural language, which have been found to improve the performance of LLMs on math reasoning tasks. Beyond in-context learning approaches, other work Gou et al. (2023); Mickus et al. (2024) fine-tuned medium-scale language models such as Code-Llama Roziere et al. (2023) on code reasoning paths and achieved comparable performance to closed-source models on math reasoning tasks. However, both in-context and fine-tuned approaches are predominantly evaluated by the correctness of the final answer, overlooking the intermediate programs and how their quality might affect the reasoning process. This work focuses on exploring these programs to accurately evaluate the problem-solving abilities of code-assisted LLMs.
Evaluation of Intermediate Steps This work also relates to several studies that evaluate the intermediate reasoning steps of LLMs in natural language (Golovneva et al., 2022; Hao et al., 2024; Jie et al., 2024; Li et al., 2024) targeting a multitude of evaluation dimensions, such as robustness and faithfulness, or error identification and correction in several reasoning tasks. These works either use a human-written reference chain to compare against the evaluated chain or employ a capable LLM to localize errors and judge the quality of the reasoning chains. Code intermediate steps are unexplored, therefore we aim to analyze the quality of these and how they affect LLMs end performance.
Evaluation of code generated by LLMs. To assess code generated by LLMs, previous work focuses on test-based evaluation and functional correctness, such as (Chen et al., 2021; Liu et al., 2024), others (Ren et al., 2020; Eghbali and Pradel, 2022; Zhou et al., 2023b) proposed metrics to measure how similar a generation is to a reference human-written code, which is expensive to get and doesn’t usually account for generation diversity. (Tong and Zhang, 2024) employed a capable LLM as a judge to localize errors in generated programs on code generation tasks. We draw inspiration from their method to evaluate generated programs, on math reasoning tasks, using an LLM-Judge. Finally, (Dou et al., 2024) analyzed characteristics of LLMs generated programs, on code generation tasks, in terms of code complexity, number of API calls, and types of errors, i.e: bugs, which cause programs to fail. Although insightful, the analysis concluded with a general type of error, commonly shared across different LLMs, namely logic error. In this work, we delve deeper to understand the logic error LLMs commit when writing code to solve a math problem.
3 Evaluation of Intermediate Programs
The main focus of this evaluation is to assess the generated programs of code-assisted LLMs beyond end-to-end evaluation and single-score metrics. We take a step further towards understanding the logic LLMs employ to solve math problems. Therefore, we evaluate their use of math concepts in generated programs. A math concept refers to the principle used to solve a specific mathematical problem, e.g: using the Euclidean Algorithm to find the Greatest Common Divisor of two numbers. Evaluating generated programs for their use of math concepts is analogous to being right for the right reason, as using the math concept in the solution ensures reasoning that is more aligned to human reasoning, and one that can be easily verified and trusted.
3.1 Evaluation Set-up
Given a math problem in natural language, we prompt an LLM to generate a Python program that solves the problem. We discard any generated programs that fail to parse 111 We resolve import errors and global misindentation (where the entire body is misindented). See Appendix A.2 for more details. from our analysis. We provide a list of the most common errors of the generated programs in Appendix D.1
3.2 Evaluation Dataset
We use math problems from two datasets, namely ASDiv Miao et al. (2020) and MATH500 Hendrycks et al. (2021); Lightman et al. (2023). These two sets include diverse problems ranging from grade-school to high-school competition-level questions. We exclude simple problems from the ASDiv data and focus only on more complex skills such as solving a system of equations, finding the greatest common divisor, or the least common multiple. The MATH500 set, on the other hand, has been reported to be more challenging for many LLMs Qwen et al. (2025); Hendrycks et al. (2021), Therefore, we use it to study how LLMs modify their generated programs to tackle more challenging math problems.
3.3 Manual Analysis
Motivated by the lack of fine-grained assessment of LLMs generated code for math reasoning tasks, we conduct a detailed human-manual analysis of the generated programs to answer our first two research questions on whether LLMs ground their generated programs to math concepts and how this grounding is implemented in the generated output. This analysis uses only a subset of the evaluation dataset. sized $300$ programs randomly sampled from the entire set, and consists of two main steps:
Sectioning the programs
We section each program into three blocks: (1) Transcription: This part of the program transcribes the information from the question into variables. (2) Processing: In this section, variables are manipulated via operations and function calls to arrive at the answer. (3) Results collection: Return the final answer of the processing section. Some sections can be fused into one line of code that represents for example processing and returning results at the same step.
Analyzing Processing Sections
We inductively analyze processing code-lines with consideration of the following aspects: logic, implementation style, coherence, verification effort, and trustworthiness.
We check if the code-lines are organized into a coherent sequence of steps that represent a valid implementation of some math concept, that a human can verify against existing possible implementations of the concept.
Additionally, we check whether all steps of a solution are explicitly generated rather than inferred by the model. Inferred steps might involve calculations models can’t compute correctly, due to their proven shortcomings in arithmetic calculations Cobbe et al. (2021); Lewkowycz et al. (2022); Gao et al. (2023), which makes verifying and trusting these solutions harder.
Most importantly, we check how math concepts are implemented in the generated programs. The grounded implementation can happen in several styles, either in terms of primitive operations only, for very simple math problems, and from-scratch implementations, or by calling functions from related math libraries that represent a more abstracted form of the solution. We also observe some implementations relying on brute-force loops that search the space of all possible answers one by one, or lack the processing section altogether.
We organize all these observed trends into six categories that make up our proposed taxonomy of grounding to math concepts and its style.
3.4 Proposed Taxonomy
Given the observations from the manual analysis, we propose to categorize LLMs’ generated programs into six mutually exclusive classes, three of which represent programs grounded to math concepts, in different implementation styles, while the other three represent ungrounded programs that mainly rely on memorized information or exhaustive searches to find the final answer.
  1. 1.
    Conceptual through library calls. Reference a math concept through calls to relevant math libraries, standard, or external.
  2. 2.
    Primitive programs: solutions are expressed in terms of the primitive operations only due to problem simplicity, where no library functionality can be called or implemented.
  3. 3.
    From-scratch Implementation of a library functionality. Instead of a call to a library function, the model implements the same functionality from scratch. The model either inlines this implementation in the generated code or write it as a custom function to be called when required.
  4. 4.
@@ 3.4 Proposed Taxonomy
    Brute-Force: The program searches through all possible values to find the answer without guiding the search with any math knowledge.
@@
    Disorganized: The program consists of incoherent steps that seem to be a mix of the previous classes. Usually include variables defined but not used, or the opposite.
@@
    No Logic: These programs skip the processing section altogether, merely returning a result without explicitly generating the steps to arrive at it. (Generating the logic as comments in natural language is also considered No Logic.)

### v2

# Evaluating Intermediate Reasoning of Code-Assisted Large Language Models for Mathematics

@@ Abstract
Assisting LLMs with code generation improved their performance on mathematical reasoning tasks. However, the evaluation of code-assisted LLMs is generally restricted to execution correctness, lacking a rigorous evaluation of their generated programs. In this work, we bridge this gap by conducting an in-depth analysis of code-assisted LLMs generated programs in response to math reasoning tasks, with a focus on evaluating the soundness of the underlying reasoning processes. For this purpose, we assess the generations of five LLMs, on several math datasets, both manually and automatically, and propose a taxonomy of generated programs based on their logical soundness. Our findings show that the capabilities of models significantly impact the logic implemented to solve the problem. Closed-source LLMs ground their programs in mathematical concepts, whereas open-source models often resort to unsound reasoning, relying on memorized information and exhaustive searches. Furthermore, increasing the difficulty of problems decreases sound generations for all models, revealing a critical shortcoming of LLMs on complex mathematics, contrary to what accuracy metrics suggest. Our work highlights the need for more holistic evaluations of code-assisted LLMs beyond execution accuracy metrics, toward a better understanding of LLMs’ limits in the math domain.
Evaluating Intermediate Reasoning of Code-Assisted Large Language Models for Mathematics
@@ 1 Introduction
Large Language Models (LLMs) have recently achieved outstanding performance on complex reasoning tasks such as mathematical reasoning, powered by scale and multi-step reasoning approaches. Particularly, the Chain-of-Thought (CoT) Wei et al. (2022) requires an LLM to generate the explicit reasoning steps, before generating the final answer. Despite its success, investigating CoT reasoning steps revealed critical flows of LLMs, such as committing calculation errors Gao et al. (2023) and generating false positive chains Lyu et al. (2023), i.e: containing reasoning errors yet generating correct final answers. Code-assisted reasoning approaches Gao et al. (2023); Chen et al. (2022); Lyu et al. (2023); Gou et al. (2023); Yue et al. (2023); Das et al. (2024) proposed to solve these problems by instructing LLMs to generate programmatic reasoning steps instead, e.g: Python programs, and delegate their execution to an external interpreter, which ensures precise calculations and faithfulness. Such approaches have been found to further improve LLMs’ performance on math tasks.
@@
However, performance improvement is predominantly measured by the correctness of the execution outcome Gao et al. (2023); Chen et al. (2022); Gou et al. (2023), rather than the quality of the generated programs and the underlying reasoning process.
This is problematic, as the generated programs can rely on exhaustive searches or memorized information to produce correct answers, leading to untrusted and more difficult-to-verify programs. Figure 1 shows programs generated by several LLMs using these hacks when solving math problems.
@@
Figure 1: Program generated by GPT-4 (left) that uses a brute force search to find the answer to the input question, and StarCoder2 (right) that depends on memorized information rather than generating solution steps to solve the input question. Both input questions are from ASDiv dataset.
@@
The goal addressed in this work is to evaluate the reasoning processes of code-assisted LLMs when solving mathematical tasks by analyzing their generated programs. In our evaluation, we focus on assessing the soundness of the logic governing LLMs’ solutions and its impact on end performance. We also assess other aspects of the generated programs, such as API calls, complexity, and the most common errors, for a more comprehensive evaluation.
@@
Our assessment begins by manually analyzing a subset of the generated programs, produced by GPT4o-mini, GPT4, Qwen2.5, Llama3, and StarCoder2, as they solve math problems from the ASDiv and MATH500 datasets (3.4). Given the observations from the manual analysis, we design a taxonomy reflecting the different logic types used by evaluated models (3.5). To extend the analysis to the complete set of generated programs, we employ two automated evaluation methods: an LLM-Judge using the latest o3-mini model, and our proposed method, Code-Structure Judge, that trains a Decision Tree classifier with features extracted from programs’ Abstract Syntax Tree (3.6). We find Code-Structure judge to outperform LLM-judge, with $81\%$ accuracy against $73\%$ (5.1); therefore, we employ it for the large-scale evaluation of all generated programs.
@@
To the best of our knowledge, this is the first work to analyze generated programs of code-assisted LLMs on math reasoning tasks. We summarize our findings from Section (5.2) below:
  • LLMs’ capabilities influence the type of reasoning they implement to tackle a math task. GPT models and Qwen frequently generate sound programs that are grounded in math concepts, while open-source LLMs resort to unsound reasoning, exploiting memorized information or brute-force searches to find final answers.
  • Difficult mathematical problems significantly decrease the distribution of sound generations, even for capable LLMs.
  • Code-assisted LLMs are not consistent in the type of logic they employ to approach problems within the same math subdomain.
  • Code-assisted LLMs can achieve comparable performance regardless of the type of logic they employ, hindering the trustworthiness of their generated programs.
Our in-depth evaluation highlights the need for more holistic assessment of LLMs’ generations, beyond accuracy metrics, that fail to reflect models’ actual capabilities and limits in the math domain.
2 Related Work
Programs as Intermediate Steps for Math Reasoning.
Code-assisted reasoning approaches such as Chen et al. (2022); Gao et al. (2023); Imani et al. (2023); Lyu et al. (2023); Das et al. (2024) prompt LLMs to generate programs instead of intermediate steps in natural language, which have been found to improve the performance of LLMs on math reasoning tasks. Beyond in-context learning approaches, other work Gou et al. (2023); Yue et al. (2023) fine-tuned medium-scale language models such as Code-Llama Roziere et al. (2023) on code reasoning paths and achieved comparable performance to closed-source models on math reasoning tasks. However, both in-context and fine-tuned approaches are predominantly evaluated by the correctness of the final answer, overlooking the intermediate programs and how they implement the reasoning process. This work focuses on exploring these programs to evaluate the problem-solving abilities of code-assisted LLMs accurately.
Evaluation of Intermediate Steps.
This work also relates to several studies that evaluate the intermediate reasoning steps of LLMs in natural language (Golovneva et al., 2022; Hao et al., 2024; Jie et al., 2024; Li et al., 2024) targeting a multitude of evaluation dimensions, such as robustness and faithfulness, or error identification and correction in several reasoning tasks. These works either use a human-written reference chain to compare against the evaluated chain or employ a capable LLM to localize errors and judge the quality of the reasoning chains. Code intermediate steps are unexplored; therefore, we aim to analyze the quality of these and how they affect LLMs’ end performance.
Evaluation of code generated by LLMs.
To assess code generated by LLMs, previous work focuses on test-based evaluation and functional correctness, such as (Chen et al., 2021; Liu et al., 2024), others (Ren et al., 2020; Eghbali and Pradel, 2022; Zhou et al., 2023b) proposed metrics to measure how similar a generation is to a reference human-written code, which is expensive to get and doesn’t usually account for generation diversity. (Tong and Zhang, 2024) employed a capable LLM as a judge to localize errors in generated programs on code generation tasks. We draw inspiration from their method to evaluate generated programs, on math reasoning tasks, using an LLM-Judge. Finally, (Dou et al., 2024) analyzed characteristics of LLMs generated programs, on code generation tasks, in terms of code complexity, number of API calls, and types of errors, i.e: bugs, which cause programs to fail. Although insightful, the analysis concluded with a general type of error, commonly shared across different LLMs, namely logic error. In this work, we delve deeper into understanding the logic errors that LLMs commit when writing code to solve a math problem.
3 Evaluation of Intermediate Programs
The primary focus of this evaluation is to assess the soundness of underlying reasoning processes, i.e: the logic implemented in the generated programs of code-assisted LLMs. We consider a program to be logically sound if it grounds the implementation in a math concept, where a math concept refers to the principle used to solve a specific mathematical problem, e.g: using the Euclidean Algorithm to find the Greatest Common Divisor of two numbers. Sound programs align more with human reasoning and can be easily verified and trusted. In contrast, unsound programs are harder to verify and can’t guarantee finding a solution to the given problem. Additionally, we examine the characteristics of generated programs in terms of cyclomatic complexity, types of errors, and API calls to provide a more comprehensive evaluation.
3.1 Evaluation Set-up
Given a math problem in natural language, we prompt an LLM to generate a Python program that solves the problem. In the initial analysis, we report the characteristics of the generated programs. Then, we discard the programs that fail to parse 111 We resolve import errors and global misindentation (where the entire body is misindented). See Appendix A.2 for more details. to evaluate logical soundness.
3.2 Evaluation Dataset
We evaluate LLMs on two popular math datasets, namely ASDiv Miao et al. (2020) and MATH500 Hendrycks et al. (2021); Lightman et al. (2023). These two sets include diverse problems ranging from grade-school to high-school competition-level questions. We exclude simple problems from the ASDiv data and focus only on more complex skills such as solving a system of equations, finding the greatest common divisor, or the least common multiple. The MATH500 set, on the other hand, has been reported to be more challenging for many LLMs Qwen et al. (2025); Hendrycks et al. (2021), Therefore, we utilize it to investigate how LLMs modify the logic in their generated programs to tackle more complex math problems.
3.3 Initial Analysis: Programs Characteristics
We analyze the generated programs in terms of their API calls to relevant math libraries, cyclomatic complexity, and the most common errors they produce. Figure 2 demonstrates API calls and cyclomatic complexity of programs generated by the evaluated models on the MATH500 problems. We observe that some models exploit symbolic computations through the use of sympy, while others prefer numerical approaches through the math library. In contrast, one LLM relies much less on external dependencies, with extremely low usage counts across the board. On the other hand, the distribution of cyclomatic complexity, in Figure 2(a), shows that generated programs by some models have high complexity values, indicating higher branching code that might be due to the complex conditional logic these LLMs are implementing to solve the problem. Finally, we present a list of the most common errors of programs generated on the MATH500 dataset in Appendix D.1.
(a) Cyclomatic Complexity
(b) API calls
Figure 2: Cyclomatic complexity and API Calls of programs generated by evaluated models in response to MATH500.
3.4 Manual Analysis: Programs Logical Soundness
We conduct a detailed human-manual analysis of the generated programs to investigate the type of logic code-assisted LLMs employ when implementing their reasoning in a Python program. This analysis considers only a subset of the evaluation dataset, sized $300$ programs randomly sampled from the entire set, and consists of two main steps:
Sectioning the programs
We section each program into three blocks: (1) Transcription: This part of the program transcribes the information from the question into variables. (2) Processing: In this section, variables are manipulated via operations and function calls to arrive at the answer. (3) Results collection: Return the final answer of the processing section. Some sections can be combined into a single line of code that represents, for example, processing and returning results simultaneously.
Analyzing Processing Sections
We inductively analyze processing code lines, considering the following aspects: logic, implementation style, coherence, verification effort, and trustworthiness. We verify that the code lines are organized into a coherent sequence of steps resembling an implementation of a mathematical concept, that a human can verify against existing implementations of the concept. Additionally, we verify that all steps of a solution are explicitly generated, rather than inferred by the model. Inferred steps might involve imprecise computations, due to LLMs’ proven shortcomings in arithmetic calculations Cobbe et al. (2021); Lewkowycz et al. (2022); Gao et al. (2023), which makes verifying and trusting these solutions harder. Finally, we check how math concepts are implemented in the generated programs. The grounded implementations can occur in several styles, including those that utilize only primitive operations for straightforward math problems, from-scratch implementations, or by calling functions from related math libraries that represent a more abstracted form of the solution. Notably, we observe other trends in some generated programs, such as relying on brute-force loops that search the space of all possible answers one by one, or lacking a processing section altogether, directly returning an answer.
3.5 Proposed Taxonomy of Programs
Given the observations from the manual analysis, we propose to categorize LLMs’ generated programs into six mutually exclusive classes, three of which represent programs with logically sound and grounded reasoning, but vary in implementation style, while the other three represent ungrounded programs, with unsound reasoning, that mainly rely on memorized information or exhaustive searches to find the final answer.
@@ 3.5 Proposed Taxonomy of Programs
    Conceptual programs through library calls. Reference a math concept through calls to relevant math libraries, standard, or external.
@@
    Primitive programs are expressed in terms of the primitive operations only due to problem simplicity, where no library functionality can be called or implemented.
@@
    From-scratch Implementation of a library functionality. Instead of a call to a library function, the model implements the same functionality from scratch. The model either inlines this implementation in the generated code or writes it as a custom function to be called when required.
@@
    Brute-Force programs that search through all possible values to find the answer without guiding the search with any math knowledge.
@@
    Disorganized programs consist of incoherent steps that seem to be a mix of the previous classes. Usually includes variables defined but not used, or the opposite.
@@
    No Logic programs skip the processing section altogether, merely returning a result without explicitly generating the steps to arrive at it. (Generating the logic as comments in natural language is also considered No Logic.)